# The Unbounded Denominators Conjecture

### v3

@@ 1.1.4. The dimension bound can be leveraged with growing level NN
The meaning of the requisite inflation rate is clarified in § 4, with Proposition 4.3.5 and Remark 4.3.8. It turns out that the logarithmically inflated holonomy rank (dimension) bound by $O(N^{3}\log{N})$ is sufficient for the desired proof by contradiction (but an $O(N^{3+1/\log{\log{N}}})$ or worse form of bound would not suffice); and this is what we ultimately prove. Getting to this degree of precision creates however some additional challenges. A straightforward elaboration of André’s original argument in [And89, Criterium VIII 1.6], taking the number of variables $d\to\infty$ and involving the $\sup_{|z|=r}\log{|\varphi|}$ growth term of loc.cit. in place of our mean (integrated) growth term $m(r,\varphi)$ in (2.0.7), leads quite easily to an $O(N^{5})$ dimension bound; and by a further work explicitly with the cusps of the Fuchsian uniformization $D(0,1)/\Gamma_{N}\cong\mathbf{C}\smallsetminus\mu_{N}$, and an appropriate Riemann map precomposition, it is possible to further reduce that down to an $O(N^{4})$. See Remark 5.2.18. This does not suffice to conclude the proof. Going further requires an intrinsic improvement into André’s dimension bound itself: the reduction of the supremum term to the integrated term in the numerator of (2.0.7).
@@ 2. The arithmetic holonomicity theorem
The novel point of the bound (2.0.3) is the integrated term in the numerator instead of a supremum term. This is critical for our proof of the unbounded denominators conjecture to have the numerator in (2.0.3), which measures the growth of $\varphi$, expressed as a Nevanlinna mean proximity function (or a characteristic function).
@@
Here by trivial local monodromy around $x=\alpha$ we mean that there exist a complex neighborhood $U_{\alpha}\ni\alpha$ and meromorphic functions $g_{1},\dots,g_{n}\in\mathcal{M}(U_{\alpha})$ on $U_{\alpha}$, where $n$ is the order of $L$, such that $g_{1},\dots,g_{n}$ form a $\mathbf{C}$-basis of the solution space of $L(f)=0$ on $U_{\alpha}\smallsetminus\{\alpha\}$. This is the case if $x=\alpha$ is not a singular point of $L$, and apart from that the typical examples singular at $x=0$ include $L_{n}=x\frac{d}{dx}-n$ for $n\in\mathbf{Z}\smallsetminus\{0\}$, of solution space $\ker{L_{n}}=\mathbf{C}\cdot x^{n}$.
@@
Let $0\in U\subset\mathbf{C}$ be an open subset containing the origin. If the uniformization radius of the pointed Riemann surface $(U,0)$ is strictly greater than $1$, then the algebra $\mathcal{V}(U,\mathbf{Z})$ is finite dimensional as a $\mathbf{Q}(x)$-vector space.
More precisely, let $p(x)\in\mathbf{Q}(x)\smallsetminus\mathbf{Q}$ be a non-constant rational function without poles in $U$, and let $\varphi(z):\overline{D(0,1)}\to U$ a holomorphic map taking $\varphi(0)=0$ with $|\varphi^{\prime}(0)|>1$. If
@@
The pulled back space $\varphi^{*}\mathcal{H}(U,x(t),\mathbf{Z})\subset\varphi^{*}\mathcal{H}(U,\mathbf{C})$ lies in the ring of formal power series fulfilling linear differential equations with analytic coefficients and no singularities with nontrivial local monodromies on the closed disc $\overline{D(0,1)}$. Hence, for any such function $f\in\mathcal{H}(U,x(t),\mathbf{Z})$, there exists a nonzero $g(x)\in\mathbf{Q}[p(x)]\smallsetminus\{0\}$ such that for any singular point $\alpha\in U$ of the linear operator $L$ in Definition 2.0.4, and for any local solution $h(x)$ in a small punctured neighborhood of $\alpha$, the product function $g(x)h(x)$ is holomorphic at $x=\alpha$. (The singularities of $L$ and thus $h(x)$ in $U$ are all at algebraic points, hence such a $g(x)$ exists.) Cauchy’s theorem then gives that $\varphi^{*}(gf)$ is a holomorphic function on $\overline{D(0,1)}$, and we conclude by Theorem 2.0.1. ∎
@@
We will see in §2.3 that Lemma 2.0.8 is in fact equivalent to Theorem 2.0.1. See however Remark 2.3.3 about a sketch of an improvement.
@@ 2.1. The auxiliary construction
Since $\varphi$ maps $(D(0,1),0)\to(\mathbf{C},0)$ with nonzero derivative, the inverse function theorem gives a positive radius $\rho>0$ such that
@@
is an analytic isomorphism from the connected component $\varphi^{-1}(D(0,\rho))_{0}$ containing 0 of $\varphi^{-1}(D(0,\rho))$.
@@
Let $d,\alpha\in\mathbf{N}$ and $\kappa\in(0,1)$ be parameters. Asymptotically in $\alpha\to\infty$ as $d$ and $\kappa$ are held fixed, there exists a nonzero $d$-variate formal function $F(\mathbf{x})$ of the form
@@
    all $a_{\mathbf{i,k}}\in\mathbf{Z}$ are rational integers bounded in absolute value by $\exp\big(\kappa C\alpha+o(\alpha)\big)$ for some constant $C\in\mathbf{R}$ depending only on the radius $\rho$ from (2.1.1) and on the degree and height of the rational function $p(x)\in\mathbf{Q}(x)$.
@@
We expand our sought-for formal function in (2.1.3) into a formal power series in $\mathbf{Q}\llbracket\mathbf{x}\rrbracket$ and solve $\binom{\alpha+d}{d}\sim\alpha^{d}/d!$ linear equations in the $(mD)^{d}$ free parameters $a_{\mathbf{i,j}}$. By the formal inverse function expansion, the integrality condition $p(x(t))\in\mathbf{Z}\llbracket t\rrbracket$ entails that $x(t)\in t+(t^{2}/M)\mathbf{Z}\llbracket t/M\rrbracket$ for some $M\in\mathbf{N}$ bounded in terms of the degree and height of the rational function $p(x)$. Therefore also the inverse series $t(x)\in x+(x^{2}/M)\mathbf{Z}\llbracket x/M\rrbracket$, and so $f_{i}(x(t))\in\mathbf{Z}\llbracket t\rrbracket$ entails $f_{i}(x)\in\mathbf{Z}\llbracket x/M\rrbracket$ for all $i=1,\ldots,m$. Furthermore, by (2.1.1), every power series $f_{i}(x)\in\mathbf{Q}\llbracket x\rrbracket$ is convergent on the archimedean disc $|x|<\rho$. The result then follows from the classical Siegel lemma [BG06, Lemma 2.9.1], with $e^{C}:=M/\rho$ and the degree parameter choice
@@
that brings in a Dirichlet exponent $\sim\kappa$ as $\alpha\to\infty$. That $F\not\equiv 0$ follows since at least one $a_{\mathbf{i,j}}\neq 0$ and the formal functions $f_{1},\ldots,f_{m}\in\mathbf{Q}\llbracket x\rrbracket$ are linearly independent over $\mathbf{Q}(p(x))$. ∎
@@
To estimate the last supremum under the asymptotic $\alpha\to\infty$ for fixed $d$ and $\kappa$, we note that the sum (2.1.3) is comprised of $(mD)^{d}=\exp\big(o(\alpha)\big)$ terms, all of which are of the form
@@
for some $k_{1},\ldots,k_{d}\in\{0,\ldots,D-1\}$, and with coefficients bounded in magnitude by $\exp\big(\kappa C\alpha+o(\alpha)\big)$.
Every such term is bounded in magnitude on $\mathbf{T}^{d}$ by
@@
$m\leq\inf_{\begin{subarray}{c}d\in\mathbf{N}\\ 0<\kappa<(\log{|\varphi^{\prime}(0)|})/|C|\end{subarray}}\Big\{\frac{\frac{d}{(d!)^{1/d}}\Big(1+\frac{1}{\kappa}\Big)^{\frac{1}{d}}\cdot\max\big\{\sup_{\mathbf{T}}\log{|h|},\,\sup_{\mathbf{T}}\log{|h\cdot\varphi^{*}p|}\big\}}{{\log{|\varphi^{\prime}(0)|-\kappa C}}}\Big\}$
@@
by Stirling’s asymptotic and the key point that the constant $C$ depends only on the datum $\{p;\varphi\}$, but not on $d$ nor $\kappa$.
The algebraicity of $f_{i}$ is a fortiori by the finite dimension bound (2.0.3) as all powers of $f_{i}$ satisfy $x^{*}f_{i}^{N}\in\mathbf{Z}[[t]]$ and $\varphi^{*}f_{i}^{N}\in\mathcal{O}(\overline{D(0,1)})$, for any $N\in\mathbf{N}$. ∎
@@
By a Blaschke product, we are reduced to the case that $g(z)\in\mathcal{O}^{\times}(D(0,1))$ is a functional unit: no zeros or poles inside the unit disc $D(0,1)$. In that case, the function $\log{|g|}$ is harmonic on $D(0,1)$, and the Poisson kernel formula—see (6.1.8) below—together with the canonical decomposition $\log=\log^{+}-\log^{-}$ into positive and negative parts gives the quotient representation
(2.3.2) $g(z)=\frac{\exp\Big(-\int_{\mathbf{T}}\log^{+}{\frac{1}{|g(w)|}}\cdot\frac{w+z}{w-z}\,\mu_{\mathrm{Haar}}(w)\Big)}{\exp\Big(-\int_{\mathbf{T}}\log^{+}{|g(w)|}\cdot\frac{w+z}{w-z}\,\mu_{\mathrm{Haar}}(w)\Big)}.$
@@
In this factorization, the top and bottom both are holomorphic functions of $z\in D(0,1)$, and they both are bounded in absolute value by $\leq 1$, because the Poisson kernel satisfies $\Re\Big(\frac{w+z}{w-z}\Big)>0$ for $1=|w|>|z|$. Furthermore, the bottom in (2.3.2) takes the value $\exp\Big(-\int_{\mathbf{T}}\log^{+}{|g|}\,\mu_{\mathrm{Haar}}\Big)$ at $z=0$.
Thus, in the functional unit case $g\in\mathcal{O}^{\times}(D(0,1))$ and on $D(0,1)$ rather than $\overline{D(0,1)}$, we conclude the proof of the lemma by selecting
$h(z):=\exp\Big(\int_{\mathbf{T}}\log^{+}{|g|}\,\mu_{\mathrm{Haar}}-\int_{\mathbf{T}}\log^{+}{|g(w)|}\cdot\frac{w+z}{w-z}\,\mu_{\mathrm{Haar}}(w)\Big).$
The proof of the general case is completed by the Blaschke product clearance to a functional unit on the open unit disc, and then by shrinking the radius a little bit. The details are, for example, in Nevanlinna [Nev70, §VII.1.4] or Goluzin [Gol69, §VII.5]. ∎
@@
On the other hand, with a bit more work based on the Law of Large Numbers for restricting all the sums $\sum_{i=1}^{d}k_{i}/D$ involved in Lemma 2.1.2 to be concentrated around the mean $d/2$ under our $d\to\infty$ asymptotic, the proof method of Lemma 2.0.8 does give the finer bound

### v4

@@ 1.1.7. The dimension bound can be leveraged with growing level NN
The meaning of the requisite inflation rate is clarified in § 4, with Proposition 4.3.5 and Remark 4.3.8. It turns out that the logarithmically inflated holonomy rank (dimension) bound by $O(N^{3}\log{N})$ is sufficient for the desired proof by contradiction (but an $O(N^{3+1/\log{\log{N}}})$ or worse form of bound would not suffice); and this is what we ultimately prove. Getting to this degree of precision creates however some additional challenges. A straightforward elaboration of André’s original argument in [And89, Criterium VIII 1.6], taking the number of variables $d\to\infty$ and involving the $\sup_{|z|=r}\log{|\varphi|}$ growth term of loc.cit. in place of our mean (integrated) growth term $m(r,\varphi)$ (see § 1.1.8 and § 6.1.1), leads quite easily to an $O(N^{5})$ dimension bound; and by a further work explicitly with the cusps of the Fuchsian uniformization $D(0,1)/\Gamma_{N}\cong\mathbf{C}\smallsetminus\mu_{N}$, and an appropriate Riemann map precomposition, it is possible to further reduce that down to an $O(N^{4})$. See Remark 5.2.19. This does not suffice to conclude the proof. Going further requires an intrinsic improvement into André’s dimension bound itself: the reduction of the supremum term to the integrated term in the numerator of (2.0.7).
@@ 1.1.8. Nevanlinna theory for Fuchsian groups
Remark 1.1.10 (Big $O$ and small $o$ notation, $\mathbf{N}$ and $\mathbf{N}_{>0}$ ).
We use big $O$ and small $o$ notation throughout in their usual way. We also use Vinogradov’s $\ll$ notation which is completely synonymous with the big $O$ notation, that is, $f\ll g$ has the same meaning as $f=O(g)$. Both of these notations mean that, with respect to some implicit variables, the inequality $f\leq Cg$ holds for all values of these variables sufficiently close to some implicit limit. We call (any suitable choice of) $C$ the implicit constant, and whenever we want to stress what either the implicit variables or implicit limits are in the notation, these are included as subscripts on either $o$, $O$, or $\ll$. We shall use $\mathbf{N}=\{0,1,2,\ldots\}$ to denote the natural numbers with zero, and $\mathbf{N}_{>0}$ to denote the positive integers.
@@ 2. The arithmetic holonomicity theorem
The novel point of the bound (2.0.3) is the integrated term in the numerator instead of a supremum term. It is critical for our proof of the unbounded denominators conjecture to have the numerator in (2.0.3), which measures the growth of $\varphi$, expressed as a Nevanlinna characteristic function (or, equivalently in the holomorphic case that we consider, a mean proximity function).
@@
Here by trivial local monodromy around $x=\alpha$ we mean that there exist a complex neighborhood $U_{\alpha}\ni\alpha$ and meromorphic functions $g_{1},\dots,g_{n}\in\mathcal{M}(U_{\alpha})$ on $U_{\alpha}$, where $n$ is the order of $L$, such that $g_{1},\dots,g_{n}$ form a $\mathbf{C}$-basis of the solution space of $L(f)=0$ on $U_{\alpha}\smallsetminus\{\alpha\}$. This is the case if $x=\alpha$ is not a singular point of $L$. An example at a singular point $x=0$ include $L_{n}=x\frac{d}{dx}-n$ for $n\in\mathbf{Z}\smallsetminus\{0\}$, of solution space $\ker{L_{n}}=\mathbf{C}\cdot x^{n}$; this is meromorphic (but not holomorphic) when $n<0$.
@@
Let $0\in U\subset\mathbf{C}$ be an open subset containing the origin. If the uniformization radius of the pointed Riemann surface $(U,0)$ is strictly greater than $1$, then the algebra $\mathcal{V}(U,\mathbf{Z})$ is finite-dimensional as a $\mathbf{Q}(x)$-vector space.
More precisely, let $p(x)\in\mathbf{Q}(x)\smallsetminus\mathbf{Q}$ be a non-constant rational function without poles in $U$, and let $\varphi(z):\overline{D(0,1)}\to U$ be a holomorphic map taking $\varphi(0)=0$ with $|\varphi^{\prime}(0)|>1$. If
@@
The pulled-back space $\varphi^{*}\mathcal{H}(U,x(t),\mathbf{Z})\subset\varphi^{*}\mathcal{H}(U,\mathbf{C})$ lies in the ring of formal power series fulfilling linear differential equations with analytic coefficients and no singularities with nontrivial local monodromies on the closed disc $\overline{D(0,1)}$. Hence, for any such function $f\in\mathcal{H}(U,x(t),\mathbf{Z})$, there exists a nonzero $g(x)\in\mathbf{Q}[p(x)]\smallsetminus\{0\}$ such that for any singular point $\alpha\in U$ of the linear operator $L$ in Definition 2.0.4, and for any local solution $h(x)$ of $L(h)=0$ in a small punctured neighborhood of $\alpha$, the product function $g(x)h(x)$ is holomorphic at $x=\alpha$. (The singularities of $L$ all occur at algebraic points.) Cauchy’s theorem then gives that $\varphi^{*}(gf)$ is a holomorphic function on $\overline{D(0,1)}$, and we conclude by Theorem 2.0.1. ∎
@@
We will find in § 2.3 that the bound in Theorem 2.0.1 is equal to the infimum of the bounds in Lemma 2.0.8 across all choices of the holomorphic multiplier function $h$. Therefore, in this form, Lemma 2.0.8 is in fact equivalent to our main Theorem 2.0.1; but it turns out convenient to approach the statement in this intermediate form. On the other hand, Remark 2.3.3 sketches a strengthened form of the lemma.
@@ 2.1. The auxiliary construction
Since $\varphi$ maps $(D(0,1),0)$ to $(\mathbf{C},0)$ with nonzero derivative, the inverse function theorem gives a positive radius $\rho>0$ such that
@@
is an analytic isomorphism from the connected component $\varphi^{-1}(D(0,\rho))_{0}$ of $\varphi^{-1}(D(0,\rho))$ which contains the element 0.
@@
Let $d,\alpha\in\mathbf{N}_{>0}$ and $\kappa\in(0,1)$ be parameters. Asymptotically in $\alpha\to\infty$ as $d$ and $\kappa$ are held fixed, there exists a nonzero $d$-variate formal function $F(\mathbf{x})$ of the form
@@
    all $a_{\mathbf{i,k}}\in\mathbf{Z}$ are integers bounded in absolute value by $\exp\big(\kappa C\alpha+o(\alpha)\big)$ for some constant $C\in\mathbf{R}$ depending only on the radius $\rho$ from (2.1.1) and on the degree and height of the rational function $p(x)\in\mathbf{Q}(x)$.
@@
We expand our sought-for formal function in (2.1.3) into a formal power series in $\mathbf{Q}\llbracket\mathbf{x}\rrbracket$ and solve $\binom{\alpha+d}{d}\sim\alpha^{d}/d!$ linear equations in the $(mD)^{d}$ free parameters $a_{\mathbf{i,j}}$. To begin with, we show that in the formal inverse function expansion, the integrality condition $p(x(t))\in\mathbf{Z}\llbracket t\rrbracket$ entails $x(t)\in t+(t^{2}/M)\mathbf{Z}\llbracket t/M\rrbracket$ with some $M\in\mathbf{N}_{>0}$ bounded in terms of the degree and height of the rational function $p(x)$.
Here are the details on the construction of $M\in\mathbf{N}_{>0}$. Set y:=b(p(x)p(0))/cy:=b(p(x)-p(0))/c with $b\in\mathbf{Z}\smallsetminus\{0\}$ and $c\in\mathbf{N}_{>0}$ chosen so that $y\in\left(x^{N}+x^{N+1}\mathbf{Q}\llbracket x\rrbracket\right)\cap\mathbf{Q}(x)$ for some $N\in\mathbf{N}_{>0}$. Formally, we have a Puiseux series branch expansion $x=x(y)\in\zeta y^{1/N}+y^{1/N}\overline{\mathbf{Q}}\llbracket y^{1/N}\rrbracket$, where $\zeta^{N}=1$. Eisenstein’s theorem [BG06, § 11.4] supplies an $M_{1}\in\mathbf{N}_{>0}$ (depending on $p(x)$) for which $x(y)\in\overline{\mathbf{Z}}\llbracket y^{1/N}/M_{1}\rrbracket$. On the other hand, the binomial expansion gives $(1+u)^{1/N}=\sum_{n=0}^{\infty}\binom{1/N}{n}u^{n}\in\mathbf{Z}\llbracket u/N^{2}\rrbracket$, by a simple denominator estimate. With our assumptions $p(x(t))\in\mathbf{Z}\llbracket t\rrbracket$ and $x(t)\in t+t^{2}\mathbf{Q}\llbracket t\rrbracket$ implying $y(t)=b(p(x(t))-p(0))/c\in c^{-1}\mathbf{Z}\llbracket t\rrbracket\cap\left(x^{N}+x^{N+1}\mathbf{Q}\llbracket x\rrbracket\right)=c^{-1}\mathbf{Z}\llbracket t\rrbracket\cap\left(t^{N}+t^{N+1}\mathbf{Q}\llbracket t\rrbracket\right)=t^{N}+c^{-1}t^{N+1}\mathbf{Z}\llbracket t\rrbracket$ and hence $\zeta y(t)^{1/N}=t\left(1+b_{1}t/c+b_{2}t^{2}/c+b_{3}t^{3}/c+\cdots\right)^{1/N}$ with $\zeta^{N}=1$ and some integers $b_{1},b_{2},\ldots\in\mathbf{Z}$, the binomial expansion gives $\zeta y(t)^{1/N}\in\mathbf{Z}\llbracket t/(cN^{2})\rrbracket$, and therefore $x(t)=x(y(t))\in\overline{\mathbf{Z}}\llbracket y(t)^{1/N}/M_{1}\rrbracket\subseteq\overline{\mathbf{Z}}\llbracket t/(cN^{2}M_{1})\rrbracket$. Coupled with $x(t)\in t+t^{2}\mathbf{Q}\llbracket t\rrbracket$, this supplies the requisite formula $x(t)\in t+(t^{2}/M)\mathbf{Z}\llbracket t/M\rrbracket$ with $M:=c^{2}N^{4}M_{1}^{2}$.
Now the inverse series also has $t(x)\in x+(x^{2}/M)\mathbf{Z}\llbracket x/M\rrbracket$, and so $f_{i}(x(t))\in\mathbf{Z}\llbracket t\rrbracket$ entails $f_{i}(x)\in\mathbf{Z}\llbracket x/M\rrbracket$ for all $i=1,\ldots,m$. Furthermore, by (2.1.1), every power series $f_{i}(x)\in\mathbf{Q}\llbracket x\rrbracket$ is convergent on the archimedean disc $|x|<\rho$. The result then follows from the classical Siegel lemma [BG06, Lemma 2.9.1], with $e^{C}:=M/\rho$ and the degree parameter choice
@@
that brings in a Dirichlet exponent $\sim\kappa$ as $\alpha\to\infty$.
Since the formal functions $f_{1},\ldots,f_{m}\in\mathbf{Q}\llbracket x\rrbracket$ are linearly independent over $\mathbf{Q}(p(x))$, an easy induction argument on the dimension $d$ shows that $\{f_{\mathbf{i}}\}_{\mathbf{i}\in\{1,\ldots,m\}^{d}}$ are linearly independent over $\mathbf{Q}(p(\mathbf{x}))$. For the step of this induction, simply note that a non-zero element $Q(x_{1},\ldots,x_{d+1})\in\mathbf{Q}(p(x_{1}),\ldots,p(x_{d+1}))\smallsetminus\{0\}$ specializes to a non-zero element $Q(\mathbf{x},c)\in\mathbf{Q}(p(\mathbf{x}))\smallsetminus\{0\}$ for all but finitely many arguments $c\in\mathbf{Q}$ under setting $x_{d+1}:=c$, and so a putative relation in the $d+1$ variables $(\mathbf{x},x_{d+1})$ specializes to a relation in the $d$ variables $\mathbf{x}=(x_{1},\ldots,x_{d})$.
At this point, having established the $\mathbf{Q}(p(\mathbf{x}))$-linear independence of the constituent functions $f_{\mathbf{i}}$, the property $F\not\equiv 0$ follows since at least one $a_{\mathbf{i,j}}\neq 0$ in the form (2.1.3). ∎
@@
To estimate the last supremum under the asymptotic $\alpha\to\infty$ for fixed $d$ and $\kappa$, we note that (2.2.1) expands from (2.1.3) into a $\mathbf{Z}$-linear combination of $(mD)^{d}=\exp\big(o(\alpha)\big)$ terms of the form
@@
for some $k_{1},\ldots,k_{d}\in\{0,\ldots,D-1\}$, and with coefficients bounded in magnitude by the quantity $\exp\big(\kappa C\alpha+o(\alpha)\big)$. Every such term is bounded in magnitude on $\mathbf{T}^{d}$ by
@@
$m\leq\inf_{\begin{subarray}{c}d\in\mathbf{N}_{>0}\\ 0<\kappa<(\log{|\varphi^{\prime}(0)|})/|C|\end{subarray}}\left\{\frac{\frac{d}{(d!)^{1/d}}\Big(1+\frac{1}{\kappa}\Big)^{\frac{1}{d}}\cdot\max\big\{\sup_{\mathbf{T}}\log{|h|},\,\sup_{\mathbf{T}}\log{|h\cdot\varphi^{*}p|}\big\}}{{\log{|\varphi^{\prime}(0)|-\kappa C}}}\right\}$
@@
by Stirling’s asymptotic and the key point that the constant $C$ depends only on $P$ and $\varphi$ but not on either $d$ or $\kappa$.
The algebraicity of $f_{i}$ follows a fortiori by the finite dimension bound (2.0.3), since all powers of $f_{i}$ satisfy $x^{*}f_{i}^{N}\in\mathbf{Z}\llbracket t\rrbracket$ and $\varphi^{*}f_{i}^{N}\in\mathcal{O}(\overline{D(0,1)})$, for any $N\in\mathbf{N}_{>0}$. ∎
@@
This is in [Nev70, § VII.1.4, Theorem on p. 187] or [Gol69, § VII.5], in the more general setting of meromorphic maps $g(z)$; with the corresponding statement replacing $\int_{\mathbf{T}}\log^{+}{|g|}\,\mu_{\mathrm{Haar}}$ by the full Nevanlinna characteristic § 6.1.1 of $g$. We present the argument for the reader’s convenience, sticking to the holomorphic case of our statement. The statement is, of course, trivial for the zero function; we assume $g\not\equiv 0$. Let $a_{1},\ldots,a_{k}$ be the finitely many zeros in the closed unit disc $\overline{D(0,1)}=\{|z|\leq 1\}$ of the nonzero meromorphic function $g:\overline{D(0,1)}\to\mathbf{C}$. (The latter, we recall, means by definition that $g$ is meromorphic on some open neighborhood of the closed disc; hence the finiteness of the set of zeros that lie in the closed disc.) Let $n_{i}\in\mathbf{N}$ be the multiplicity of the zero $a_{i}$. The Blaschke product
$B(z):=\prod_{i=1}^{k}\left(\frac{z-a_{i}}{1-\overline{a_{i}}z}\right)^{n_{i}}\,:\,D(0,1)\to D(0,1)$
is a holomorphic self-map $D(0,1)\to D(0,1)$ of the unit disc that preserves its boundary, and
$G(z):=\frac{g(z)}{B(z)}\in\mathcal{O}^{\times}(D(0,1))$
@@
is a functional unit on the open disc: a nowhere vanishing holomorphic function on $D(0,1)$. The function $\log{|G|}:D(0,1)\to\mathbf{R}$ is therefore harmonic, and so the Poisson kernel formula — see (6.1.10) below for a review — together with the canonical decomposition $\log=\log^{+}-\log^{-}$ into positive and negative parts gives the quotient representation
$G(z)=\frac{\exp\Big(-\int_{|z|=r}\log^{+}{\frac{1}{|G(w)|}}\cdot\frac{w+z}{w-z}\,\mu_{\mathrm{Haar}}(w)\Big)}{\exp\Big(-\int_{|z|=r}\log^{+}{|G(w)|}\cdot\frac{w+z}{w-z}\,\mu_{\mathrm{Haar}}(w)\Big)}$
on a neighborhood of the closed disc $|z|\leq r$, for every $r<1$. Therefore, taking $r\to 1^{-}$, we have quotient representation on the open disc $D(0,1)$:
(2.3.2) $G(z)=\frac{\exp\Big(-\int_{\mathbf{T}}\log^{+}{\frac{1}{|G(w)|}}\cdot\frac{w+z}{w-z}\,\mu_{\mathrm{Haar}}(w)\Big)}{\exp\Big(-\int_{\mathbf{T}}\log^{+}{|G(w)|}\cdot\frac{w+z}{w-z}\,\mu_{\mathrm{Haar}}(w)\Big)}=\frac{\exp\Big(-\int_{\mathbf{T}}\log^{+}{\frac{1}{|g(w)|}}\cdot\frac{w+z}{w-z}\,\mu_{\mathrm{Haar}}(w)\Big)}{\exp\Big(-\int_{\mathbf{T}}\log^{+}{|g(w)|}\cdot\frac{w+z}{w-z}\,\mu_{\mathrm{Haar}}(w)\Big)}.$
@@
In this factorization, the top and bottom both are holomorphic functions of $z\in D(0,1)$, and they both are bounded in absolute value by $\leq 1$, because the Poisson kernel satisfies $\Re\Big(\frac{w+z}{w-z}\Big)>0$ for $1=|w|>|z|$. Furthermore, the bottom in (2.3.2) takes the value $\exp\Big(-\int_{\mathbf{T}}\log^{+}{|g|}\,\mu_{\mathrm{Haar}}\Big)$ at $z=0$. Therefore, on the open disc $D(0,1)$, the definition
$h(z):=\exp\left(\int_{\mathbf{T}}\log^{+}{|g|}\,\mu_{\mathrm{Haar}}-\int_{\mathbf{T}}\log^{+}{|g(w)|}\cdot\frac{w+z}{w-z}\,\mu_{\mathrm{Haar}}(w)\right)$
fulfills $h(0)=1$ and $\sup_{D(0,1)}|h|\leq\int_{\mathbf{T}}\log^{+}{|g|}\,\mu_{\mathrm{Haar}}$, but then also (2.3.2) taken with $g=BG$ gives
$\quad\sup_{D(0,1)}\left\{\log{|hg|}\right\}=\sup_{D(0,1)}\left\{\log{|hBG|}\right\}\leq\sup_{D(0,1)}\left\{\log{|hG|}\right\}$
$=\sup_{|w|<1}\left\{\int_{\mathbf{T}}\log^{+}{|g|}\,\mu_{\mathrm{Haar}}-\int_{\mathbf{T}}\log^{+}{\frac{1}{|g(w)|}}\cdot\frac{w+z}{w-z}\,\mu_{\mathrm{Haar}}(w)\right\}$
$\leq\int_{\mathbf{T}}\log^{+}{|g|}\,\mu_{\mathrm{Haar}}.$
This constructs the desired quotient representation except on the open disc $D(0,1)$ rather than on a neighborhood of the closed disc. The full statement bootstraps from this by the following limiting argument; this is where the $\varepsilon>0$ emerges in the statement of the theorem. Taking a small enough $\delta>0$ (to be chosen at the end in dependence on $\varepsilon$) such that $g$ is a holomorphic function on $\overline{D(0,1+\delta)}$, we apply the preceding to the holomorphic function $\widetilde{g}(z):=g((1+\delta)z)$ on $\overline{D(0,1)}$. We obtain a holomorphic function $\widetilde{h}:D(0,1)\to\mathbf{C}$ such that $\widetilde{h}(0)=1$ and
$\max\left\{\sup_{D(0,1)}\log|\widetilde{h}|,\sup_{D(0,1)}\log|\widetilde{h}\widetilde{g}|\right\}\leq\int_{\mathbf{T}}\log^{+}{|\widetilde{g}|}\,\mu_{\mathrm{Haar}}.$
Define $h(z):=\widetilde{h}(z/(1+\delta))$, a holomorphic function on $D(0,1+\delta)\supset\overline{D(0,1)}$ with $h(0)=1$. Then
$\max\big\{\sup_{\mathbf{T}}\log{|h|},\,\sup_{\mathbf{T}}\log{|hg|}\big\}\leq\int_{\mathbf{T}}\log^{+}{|\widetilde{g}|}\,\mu_{\mathrm{Haar}}=\int_{|z|=1+\delta}\log^{+}{|g(z)|}\,\mu_{\mathrm{Haar}}.$
We get what we want upon choosing $\delta=\delta(\varepsilon)>0$ small enough to have $\int_{|z|=1+\delta}\log^{+}{|g(z)|}\,\mu_{\mathrm{Haar}}\leq\int_{\mathbf{T}}\log^{+}{|g|}\,\mu_{\mathrm{Haar}}+\varepsilon$. ∎
@@
On the other hand, with a bit more work based on the Law of Large Numbers, we could restrict the auxiliary construction (2.1.3) to only admit those exponent vectors $\mathbf{k}\in\{0,\ldots,D-1\}^{d}$ that have $\sum_{i=1}^{d}k_{i}/D$ concentrated around the expectation $d/2$. With such a variant of Lemma 2.1.2, the same argument leads to the finer bound